# Overview of AI-Debater 2023: The Challenges of Argument Generation Tasks

### v1

@@ 1 Introduction
Argument and debate are fundamental capabilities of human intelligence, essential for a wide range of human activities, and common to all human societies. Argumentation [5; 16; 27] takes the human logical argumentation process as the research object, and is a research field involving logic, philosophy, language, rhetoric, computer science and education. Striving to enable models to automatically understand and generate argument texts, computational argumentation, a newly emerging research field, is obtaining increasing attention from the research community [23]. Depending on the task objectives, computational argumentation tasks can be divided into two aspects, argument mining and argument generation.
With the rapid development of modern technology, online forums like ChangeMyView allow people to freely exchange opinions on specific topics, making them suitable data sources for argument generation tasks, especially for designing artificial debaters, as online forums closely resemble real-world debates. Initial research in this field has focused on analyzing ChangeMyView data [21; 30] to summarize the key factors of persuasive arguments.
For an extended period, the field of argument mining has been particularly active. Li et al.[li2022structure] proposed a Structure-Aware Argument Encoder (SAE) in their work, which enhances the ability to capture structural information in the analysis of scientific literature discourse by distinguishing between framing words and topic words in sentences and incorporating paragraph-level positional information. Additionally, some researchers have integrated knowledge graph structures into argument mining tasks. Yuan et al.[yuan2021leveraging] constructed a knowledge graph for external knowledge, improving the model’s ability to identify interactive argument pairs. Liang et al.[liang2023hi] proposed a hierarchical argumentation graph structure and introduced a text-graph multi-modal pre-training framework.
Recently, large language models, such as OpenAI ChatGPT and GPT-4 [7], PaLM [8], and LlaMAs [25; 26] have achieved great success and demonstrated remarkable performance in text generation tasks. Therefore, to align the field of computational argumentation with the development trend of large language models, we have organized the AI-Debater 2023 Challenge [^2] This event is an CCAC 2023 task sponsored by Fudan University.. This challenge focuses on generation tasks, including two tracks: counter-argument generation (Track 1) and claim-based argument generation (Track 2). In Track 1, we introduce the task of generating counter-argument based on given topic; while in Track 2, we introduce the task of generating argument based on given claim. We provide two datasets in this task, one for each track.
@@ 2.1 Counter-Argument Generation
Datasets for counter-argument generation mainly establish the rebuttal relationship in the conversation using automatic methods such as citation or reply detection [13; 11]. Seaman [21] proposed CMV dataset, including the citation relationship between original posts and their corresponding replies. Bolton et al. [6] introduced Kialo, a dataset for sentence-level argument stance classification, which can also be applied to counter-argument generation task. Lin et al. [15] introduced ArgTersely, a dataset for sentence-level counter-argument generation, this dataset is obtained by manual annotation.
Early work [11; 12] focus on how to introduce external knowledge into the system; Alshomary et al. [4] developed a system to identify weak points in arguments; Schiller et al. [20] developed a controlled argument generation system, which is able to generate arguments based on given information; Alshomary and Wachsmuth [2] completed it through multi-task and multi-step reasoning. Lin et al. [15] constructed argumentation instructions, and fine-tuned a large language model for this task.

### v2

@@ 1 Introduction
Argument and debate are fundamental capabilities of human intelligence, essential for a wide range of human activities, and common to all human societies. Argumentation [1, 2, 3] takes the human logical argumentation process as the research object, and is a research field involving logic, philosophy, language, rhetoric, computer science and education. Striving to enable models to automatically understand and generate argument texts, computational argumentation, a newly emerging research field, is obtaining increasing attention from the research community [4]. Depending on the task objectives, computational argumentation tasks can be divided into two aspects, argument mining and argument generation.
With the rapid development of modern technology, online forums like ChangeMyView allow people to freely exchange opinions on specific topics, making them suitable data sources for argument generation tasks, especially for designing artificial debaters, as online forums closely resemble real-world debates. Initial research in this field has focused on analyzing ChangeMyView data [5, 6] to summarize the key factors of persuasive arguments.
For an extended period, the field of argument mining has been particularly active. Li et al. [7] proposed a Structure-Aware Argument Encoder (SAE) in their work, which enhances the ability to capture structural information in the analysis of scientific literature discourse by distinguishing between framing words and topic words in sentences and incorporating paragraph-level positional information. Additionally, some researchers have integrated knowledge graph structures into argument mining tasks. Yuan et al. [8] constructed a knowledge graph for external knowledge, improving the model’s ability to identify interactive argument pairs. Liang et al. [9] proposed a hierarchical argumentation graph structure and introduced a text-graph multi-modal pre-training framework.
Recently, large language models, such as OpenAI ChatGPT and GPT-4 [10], PaLM [11], and LlaMAs [12, 13] have achieved great success and demonstrated remarkable performance in text generation tasks. Therefore, to align the field of computational argumentation with the development trend of large language models, we have organized the AI-Debater 2023 Challenge [^2] This event is an CCAC 2023 task sponsored by Fudan University.. This challenge focuses on generation tasks, including two tracks: counter-argument generation (Track 1) and claim-based argument generation (Track 2). In Track 1, we introduce the task of generating counter-argument based on given topic; while in Track 2, we introduce the task of generating argument based on given claim. We provide two datasets in this task, one for each track.
@@ 2.1 Counter-Argument Generation
Datasets for counter-argument generation mainly establish the rebuttal relationship in the conversation using automatic methods such as citation or reply detection [14, 15]. Seaman et al. [5] proposed CMV dataset, including the citation relationship between original posts and their corresponding replies. Bolton et al. [16] introduced Kialo, a dataset for sentence-level argument stance classification, which can also be applied to counter-argument generation task. Lin et al. [17] introduced ArgTersely, a dataset for sentence-level counter-argument generation, this dataset is obtained by manual annotation.
Early work [15, 18] focus on how to introduce external knowledge into the system; Alshomary et al. [19] developed a system to identify weak points in arguments; Schiller et al. [20] developed a controlled argument generation system, which is able to generate arguments based on given information; Alshomary et al. [21] completed it through multi-task and multi-step reasoning. Lin et al. [17] constructed argumentation instructions, and fine-tuned a large language model for this task.